### v1

# GR-Gaussian: Graph-Based Radiative Gaussian Splatting for Sparse-View CT Reconstruction

Abstract
3D Gaussian Splatting (3DGS) has emerged as a promising approach for CT reconstruction. However, existing methods rely on the average gradient magnitude of points within the view, often leading to severe needle-like artifacts under sparse-view conditions. To address this challenge, we propose GR-Gaussian, a graph-based 3D Gaussian Splatting framework that suppresses needle-like artifacts and improves reconstruction accuracy under sparse-view conditions. Our framework introduces two key innovations: (1) a Denoised Point Cloud Initialization Strategy that reduces initialization errors and accelerates convergence; and (2) a Pixel-Graph-Aware Gradient Strategy that refines gradient computation using graph-based density differences, improving splitting accuracy and density representation. Experiments on X-3D and real-world datasets validate the effectiveness of GR-Gaussian, achieving PSNR improvements of 0.67 dB and 0.92 dB, and SSIM gains of 0.011 and 0.021. These results highlight the applicability of GR-Gaussian for accurate CT reconstruction under challenging sparse-view conditions.
Figure 1: Results. We compare our method to two NeRF-based methods NAF(Zha, Zhang, and Li 2022), SAX-NeRF(Cai et al. 2024b)) and $R^{2}$-GS(Zha et al. 2024) in terms of visual quality and PSNR (dB). Our method mitigates needle-like artifacts and achieves superior CT reconstruction quality under sparse-view conditions.
1 Introduction
Computed Tomography (CT) is a widely used non-invasive imaging technique applied in fields such as medicine, biology, security, and industry. It utilizes X-rays to capture multi-angle 2D projections, which measure the attenuation of rays as they pass through an object(Kak and Slaney 2001). These projections are then processed to reconstruct the 3D density distribution of the object, enabling detailed internal visualization.
In many applications, tomography data acquisition is limited by the number of projections or non-uniform angular sampling. For instance, medical CT reduces the number of projections to minimize X-ray exposure, while electron tomography faces the missing wedge problem due to restricted rotation angles. These constraints make the reconstruction problem highly ill-posed. Traditional analytical methods, such as FBP and FDK(Feldkamp, Davis, and Kress 1984), are computationally efficient but often fail under these conditions, producing artifacts. Model-based optimization approaches(Andersen and Kak 1984; Ma et al. 2024b), which use nonlinear regularizers and iterative solvers (e.g., ART), achieve better results but are computationally expensive.
Deep learning-based methods(Lin et al. 2023; Chung et al. 2023; Liu et al. 2023) demonstrate strong performance but require large labeled datasets and long training times(Ma et al. 2025a, c). They also struggle with out-of-distribution objects. NeRF-based(Mildenhall et al. 2021; Ma et al. 2025b, 2024a) approaches show promise in per-case reconstruction but are extremely time-consuming due to extensive point sampling during volume rendering(Zha, Zhang, and Li 2022; Cai et al. 2024b; Zang et al. 2021). These methods highlight the trade-off between reconstruction quality and computational efficiency.
Recently, 3D Gaussian Splatting (3DGS)(Kerbl et al. 2023; Yu et al. 2024; Guédon and Lepetit 2024) has demonstrated superior quality and efficiency over NeRF in view synthesis tasks(Lu et al. 2024; Liang et al. 2024). However, when applied to sparse-view tomographic reconstruction, 3DGS-based methods exhibit characteristic needle-like artifacts. Through analysis, we identify that these artifacts arise from the retention of large Gaussian kernels with small gradients during the reconstruction process. This limitation stems from the lack of consideration for inter-point relationships, which results in biased density distributions and insufficient splitting of Gaussian kernels.
To address these issues, we propose a novel framework GR-Gaussian, which introduces graph structures to model relationships between neighboring points. Specifically, we first introduce a Denoised Point Cloud Initialization Strategy(De-Init) based on an enhanced FDK method, which reduces noise and artifacts while preserving structural details. Then, we develop a Pixel-Graph-Aware Gradient Strategy(PGA) that reconstructs the gradient computation process by incorporating inter-point information. This strategy enables more accurate splitting decisions for Gaussian kernels, effectively mitigating needle-like artifacts and enhancing spatial coherence. Finally, we design an efficient optimization strategy tailored for GR-Gaussian, ensuring robust and accurate reconstruction.
We validate our framework through extensive experiments on both simulated X-3D dataset(Zha et al. 2024) and Real-world CT datasets (Society 2024), comparing it against multiple baseline methods. The results demonstrate that GR-Gaussian outperforms existing approaches in reconstruction quality, achieving PSNR gains of 0.67 dB and 0.92 dB, and SSIM improvements of 0.011 and 0.021. Our main contributions are summarized as follows:
  • We propose a novel framework GR-Gaussian, which leverages graph-structured relationships to model objects, providing a solution for sparse-view tomographic reconstruction.
  • To address needle-like artifacts in 3DGS-based methods under sparse-view conditions, we introduce the De-Init to enhance the quality of initialization. Then we propose the PGA to improve the gradient computation process.
  • Our framework achieves state-of-the-art performance on the X-3D and real-world datasets, demonstrating its effectiveness under challenging sparse-view conditions.
2 Related Work
2.1 Tomographic Reconstruction
Computed Tomography (CT) is a critical imaging technique widely used in medicine(Hounsfield 1980; Katsuragawa and Doi 2007) and industry(De Chiffre et al. 2014; Zang et al. 2019). Sparse-view CT reconstruction, where the number of projections is limited, poses significant challenges for traditional methods. Traditional analytical methods, such as Filtered Back Projection (FBP) and its 3D counterpart FDK(Feldkamp, Davis, and Kress 1984), are computationally efficient but prone to severe streak artifacts in sparse-view scenarios, leading to degraded image quality. Iterative methods(Andersen and Kak 1984; Sidky and Pan 2008; Manglos et al. 1995; Sauer and Bouman 2002), which optimize reconstruction through regularized energy minimization, can suppress artifacts but are computationally expensive and risk losing fine structural details. Deep learning methods(Lin et al. 2023; Chung et al. 2023; Liu et al. 2023) have shown promise in CT reconstruction, particularly in sparse-view scenarios. Supervised approaches leverage large datasets to learn semantic priors for tasks like projection inpainting(Anirudh et al. 2018; Ghani and Karl 2018) and volume denoising(Chung et al. 2023; Lee et al. 2023; Liu et al. 2023, 2020), but struggle with generalization to unseen data. Self-supervised methods, inspired by NeRF(Mildenhall et al. 2021), optimize density fields using photometric losses(Cai et al. 2024b; Zha, Zhang, and Li 2022; Zang et al. 2021) but are computationally prohibitive due to extensive point sampling during rendering.
2.2 3D Gaussian Splatting
3D Gaussian Splatting has demonstrated remarkable performance in RGB tasks, such as surface reconstruction(Li et al. 2025; Lin et al. 2025; Hou et al. 2025), dynamic scene modeling(Zhu et al. 2025; Huang et al. 2025; Wu et al. 2024), human avatar creation(Qu et al. 2025; Li et al. 2024), and 3D generation, outperforming NeRF in terms of speed by leveraging highly parallelized rasterization for image rendering. Recent studies have attempted to extend 3DGS to X-ray imaging. For example, Cai(Cai et al. 2024a) empirically modified 3DGS to synthesize novel-view X-ray projections, primarily using it as a data augmentation tool for traditional algorithms rather than directly generating CT models. Zha(Zha et al. 2024) represented the density field with customized Gaussian kernels, yet they replaced the efficient rasterization with existing CT simulators. While effective in certain cases, their approach overfits sparse-view projections, leading to reconstruction failures.
Figure 2: Training pipeline of GR-Gaussian. (a) Overall training pipeline. (b) Denoised Point Cloud Initialization Strategy. (c) Pixel-Graph-Aware Gradient Strategy.
3 Preliminary
3.1 Radiative Gaussians
To represent the target object, previous approach(Zha et al. 2024) utilize a set of learnable 3D Gaussian kernels, denoted as $\mathbb{G}^{3}=\{G_{i}^{3}\}_{i=1,\cdots,M}$, which we refer to as Radiative Gaussians. Each kernel $G_{i}^{3}$ defines a local Gaussian-shaped density field, expressed as:
$G_{i}^{3}(\mathbf{x}\mid\rho_{i},\mathbf{p}_{i},\Sigma_{i})=\rho_{i}\cdot e^{\left(-\frac{1}{2}(\mathbf{x}-\mathbf{p}_{i})^{\top}\Sigma_{i}^{-1}(\mathbf{x}-\mathbf{p}_{i})\right)}$ (1)
Here, $\rho_{i}$ represents the central density, $\mathbf{p}_{i}\in\mathbb{R}^{3}$ denotes the position, and $\Sigma_{i}\in\mathbb{R}^{3\times 3}$ is the covariance matrix, all of which are learnable parameters. For optimization, $\Sigma_{i}$ is decomposed as $\Sigma_{i}=\mathbf{R}_{i}\mathbf{S}_{i}\mathbf{S}_{i}^{\top}\mathbf{R}_{i}^{\top}$, where $\mathbf{R}_{i}$ and $\mathbf{S}_{i}$ are the rotation and scale matrices, respectively. The overall density at a given position $\mathbf{x}\in\mathbb{R}^{3}$ is computed as the sum of the contributions from all kernels:
$\sigma(\mathbf{x})=\sum_{i=1}^{M}G_{i}^{3}(\mathbf{x}\mid\rho_{i},\mathbf{p}_{i},\Sigma_{i})$ (2)
Notably, this formulation omits view-dependent color since X-ray attenuation depends solely on isotropic density, rendering it ideal for CT reconstruction.
3.2 Pixel-Aware Gradient
In 3D Gaussian Splatting, the decision to split or clone a Gaussian kernel $G_{i}^{3}$ is based on the average gradient magnitude of its Normalized Device Coordinates (NDC) across multiple viewpoints. For viewpoint $v$, the NDC coordinates are $\mu_{\mathrm{ndc}}^{i,v}=(\mu_{\mathrm{ndc,x}}^{i,v},\mu_{\mathrm{ndc,y}}^{i,v})$, and the corresponding loss is $L_{v}$. An Adaptive Density Control (ADC) mechanism adjusts kernel density every 100 iterations. A splitting or cloning operation is triggered if the average gradient magnitude across $N^{i}$ viewpoints exceeds a threshold $\tau_{\mathrm{pos}}$:
$\frac{1}{N^{i}}\sum_{v=1}^{N^{i}}\left\|\frac{\partial L_{v}}{\partial\mu_{\mathrm{ndc}}^{i,v}}\right\|_{2}>\tau_{\mathrm{pos}},$ (3)
During photometric reconstruction tasks, the gradient flow through color channels drives density field optimization. For a Gaussian kernel $G_{i}^{3}$ at viewpoint $v$, the gradient contribution is computed as:
$(\mathbf{g}_{i}^{r})^{v}=\frac{\partial L_{v}}{\partial\mu_{\mathrm{ndc}}^{i,v}}=\sum_{pix=1}^{m_{v}^{i}}\sum_{j=1}^{3}\frac{\partial L_{v}}{\partial c_{j}^{pix}}\cdot\frac{\partial c_{j}^{pix}}{\partial\alpha_{v,pix}^{i}}\cdot\frac{\partial\alpha_{v,pix}^{i}}{\partial\mu_{\mathrm{ndc}}^{i,v}},$ (4)
Here, $\alpha_{v,pix}^{i}$ represents the contribution factor of a pixel to $G_{i}^{3}$, $c_{j}^{pix}$ is the intensity of the $j$-th color channel, and $m_{v}^{i}$ is the total number of pixels involved. $\alpha_{v,pix}^{i}$ is defined as a function of the Euclidean distance between the Gaussian’s center and the pixel center, decaying exponentially with increasing distance.
Within CT reconstruction tasks, where images are grayscale, gradient computation is simplified by removing the summation over color channels. The gradient contribution of a pixel under viewpoint $v$ to Gaussian $G_{i}^{3}$ is:
$(\mathbf{g}_{i}^{c^{\prime}})^{v}=\sum_{pix=1}^{m_{v}^{i}}\frac{\partial L_{v}}{\partial\alpha_{v,pix}^{i}}\cdot\frac{\partial\alpha_{v,pix}^{i}}{\partial\mu_{\mathrm{ndc}}^{i,v}}.$ (5)
This adaptation focuses on density contributions and ignores color information. The exponential decay of $\alpha_{v,pix}^{i}$ emphasizes the influence of pixels closer to the Gaussian center, enabling precise gradient updates.
4 Method
In this section, we present GR-Gaussian, a novel framework that represents objects through graph-structured relationships (Sec. 4.1). Specifically optimized for sparse-view CT reconstruction, our framework effectively tackles critical issues, most notably the emergence of needle-like artifacts (Sec. 4.2). We first introduce a Denoised Point Cloud Initialization Strategy based on an enhanced FDK method. Subsequently, a Pixel-Graph-Aware Gradient Strategy is proposed to refine the gradient computation process. Finally, we detail an optimization strategy specifically tailored for GR-Gaussian.
Figure 3: The scanned object is represented as graph-based radiative Gaussians, optimized using real X-ray projections to retrieve the density volume via voxelization.
4.1 Representing Objects with Graph-based Radiative Gaussians
As shown in Figure 3, an object is represented as a learnable 3D Gaussian graph $\mathcal{G}$, defined as:
$\mathcal{G}=(\mathcal{V},\mathcal{E},\{G_{i}^{3}(\rho_{i},\mathbf{p}_{i},\Sigma_{i})\}_{i=1}^{M}),$ (6)
where $\mathcal{V}=\{\mathbf{p}_{i}\}_{i=1}^{M}$ denotes the positions of the Gaussian kernels, $\mathcal{E}$ represents the edges connecting the kernels, and $G_{i}^{3}$ is a Gaussian kernel parameterized by density $\rho_{i}$, position $\mathbf{p}_{i}\in\mathbb{R}^{3}$, and covariance $\Sigma_{i}\in\mathbb{R}^{3\times 3}$.
The graph structure $\mathcal{G}$ is constructed to capture spatial relationships between Gaussian kernels. The vertices $\mathcal{V}$ represent kernel positions, while the edges $\mathcal{E}$ are determined using the KNN algorithm:
$\text{KNN}(\mathbf{p}_{i})=\{\mathbf{p}_{j}\mid j\in\text{argmin}_{j}\,d(\mathbf{p}_{i},\mathbf{p}_{j}),\,j=1,\cdots,K\},$ (7)
where the Euclidean distance $d(\mathbf{p}_{i},\mathbf{p}_{j})$ is given by:
$d(\mathbf{p}_{i},\mathbf{p}_{j})=\|\mathbf{p}_{i}-\mathbf{p}_{j}\|_{2}.$ (8)
To ensure bidirectional connectivity, the edge set $\mathcal{E}$ is defined as:
=$\mathcal{E}=\{(\mathbf{p}_{i},\mathbf{p}_{j})\mid\mathbf{p}_{j}\in\text{KNN}(\mathbf{p}_{i})\text{ and }\mathbf{p}_{i}\in\text{KNN}(\mathbf{p}_{j})\}.$ (9)
Each Gaussian kernel $G_{i}^{3}$ defines a local density field by combining its own density with contributions from neighboring kernels:
$G_{i}^{3}(\mathbf{x})=\rho_{i}\cdot G(\mathbf{x}\mid\mathbf{p}_{i},\Sigma_{i})+\sum_{j\in\mathcal{N}(i)}w_{ij}\cdot\rho_{j}\cdot G(\mathbf{x}\mid\mathbf{p}_{j},\Sigma_{j}),$ (10)
where $G(\mathbf{x}\mid\mathbf{p},\Sigma)$ is the Gaussian function:
$G(\mathbf{x}\mid\mathbf{p},\Sigma)=\exp\left(-\frac{1}{2}(\mathbf{x}-\mathbf{p})^{\top}\Sigma^{-1}(\mathbf{x}-\mathbf{p})\right),$ (11)
The weight $w_{ij}$ is introduced to quantify the influence of neighboring kernels based on their spatial proximity. It is defined as:
$w_{ij}=\exp\left(-\frac{\|\mathbf{p}_{i}-\mathbf{p}_{j}\|^{2}}{k}\right),$ (12)
where $\|\mathbf{p}_{i}-\mathbf{p}_{j}\|^{2}$ represents the squared Euclidean distance between kernel positions $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$, and $k$ is a scaling factor that controls the sensitivity of the weight to distance. The combined density $D(\mathbf{x})$ at a position $\mathbf{x}$ is computed as:
$D(\mathbf{x})=\sum_{i=1}^{M}G_{i}^{3}(\mathbf{x}).$ (13)
This graph-based representation is specifically designed to model the relationships between neighboring points, enabling the incorporation of inter-point information. By leveraging these relationships, the framework facilitates the optimization of update strategies, improving the accuracy of the reconstruction process.
4.2 Training Graph-based Radiative Gaussians
Our training pipeline, illustrated in Figure 2, begins with the initialization of Graph-based Radiative Gaussians from a modified FDK volume. A graph is then constructed using the KNN method to capture spatial relationships between Gaussian kernels. Projections are rasterized to compute photometric losses, while tiny density volumes are voxelized for 3D regularization. Finally, a modified Pixel-Graph-Aware Gradient Strategy is applied to densify the Gaussians.
Denoised Point Cloud Initialization Strategy
Previous methods utilize FDK to generate low-quality volumes for initialization. However, under sparse-view conditions, FDK reconstructions suffer from significant noise and artifacts, which degrade subsequent processing. To address this, as shown in Figure 2 (b), we propose a Denoised Point Cloud Initialization Strategy leveraging 3DGS characteristics. Gaussian filtering is applied to the FDK-reconstructed point cloud to suppress noise and artifacts while preserving structural details, ensuring high-quality initialization for robust and accurate reconstruction. The Gaussian filtering process is defined as:
$f^{\prime}(i)=\frac{1}{Z}\sum_{j\in\mathcal{N}(i)}G(j;\sigma_{d})\cdot f(i+j),.$ (14)
Here, $i$ denotes the coordinates of the current voxel, and $j$ represents the offset within its neighborhood $\mathcal{N}(i)$, defined as $j_{k}\in[-r,r]$ for all dimensions $k$. The Gaussian kernel is given by:
$G(j;\sigma_{d})=\exp\left(-\frac{|j|^{2}}{2\sigma_{d}^{2}}\right),$
where $|j|^{2}=j_{1}^{2}+j_{2}^{2}+\cdots+j_{n}^{2}$, and $Z=\sum_{j\in\mathcal{N}(i)}G(j;\sigma_{d})$ is the normalization factor ensuring the kernel weights sum to 1. Empty regions are excluded using a density threshold $\tau$, and $M$ points are randomly sampled as kernel positions. Following (Zha et al. 2024), Gaussian scales are set based on nearest neighbor distances, assuming no rotation. Central densities are directly queried from the denoised volume.
Pixel-Graph-Aware Gradient Strategy
Previous methods compute gradients based solely on average values, limiting the contribution of large kernels with small gradients to the densification process. To address this limitation, we propose a Pixel-Graph-Aware Gradient Strategy that integrates graph-based relationships to refine gradient computation and enhance the splitting process. This strategy leverages the density differences between Gaussian kernels and their neighbors, consistent with the characteristics of X-ray CT imaging, where similar tissues or materials exhibit approximately constant attenuation coefficients. The density difference between a Gaussian kernel $i$ and its neighbors $\mathcal{N}(i)$ is defined as:
$\Delta\rho_{ij}=|\rho_{i}-\rho_{j}|,\quad\forall j\in\mathcal{N}(i).$ (15)
The density difference $\Delta\rho_{ij}$ between a Gaussian kernel $G_{i}^{3}$ and its neighbor $G_{j}^{3}$ reflects significant variations, with larger values contributing to higher gradient magnitudes. To enhance gradient computation, we incorporate these density differences into the gradient formula. For a Gaussian kernel $G_{i}^{3}$, the augmented gradient is defined as:
$(\mathbf{g}_{i}^{c})^{v}=\sum_{pix=1}^{m_{v}^{i}}\frac{\partial L_{v}}{\partial\alpha_{v,pix}^{i}}\cdot\frac{\partial\alpha_{v,pix}^{i}}{\partial\mu_{\mathrm{ndc}}^{i,v}}+\lambda_{g}\cdot\frac{\sum_{j\in\mathcal{N}(i)}\Delta\rho_{ij}}{k},$ (16)
where $(\mathbf{g}_{i}^{c})^{v}$ denotes the augmented gradient under viewpoint $v$, $\lambda_{g}$ is a regularization parameter controlling the influence of neighboring kernels, and $\mathcal{N}(i)$ represents the set of neighbors connected to $G_{i}^{3}$ in the graph. By leveraging graph-based relationships, the augmented gradient effectively captures density variations:
$(\mathbf{g}_{i}^{c})^{v}\propto\Delta\rho_{ij},\quad\forall j\in\mathcal{N}(i).$ (17)
This approach enhances gradient estimation for large kernels, enabling more effective splitting. The increased gradient magnitude improves the representation of the underlying density field, governed by the splitting condition $\|(\mathbf{g}_{i}^{c})^{v}\|>\tau_{\mathrm{pos}}$, where $\tau_{\mathrm{pos}}$ is a predefined threshold. The Pixel-Graph-Aware Gradient Strategy integrates spatial relationships into gradient computation for 3D Gaussian Splatting, leveraging graph-based density differences to refine the splitting strategy.
Figure 4: The Pixel-Graph-Aware Gradient Strategy leverages density differences between Gaussian kernels by constructing a graph to encode point-to-point relationships, enhancing gradient computation and enabling effective splitting of large kernels with low gradients.
Density Voxelizer
We employ a voxelizer, denoted as $\mathbf{V}$, to efficiently extract a density volume $V\in R^{X\times Y\times Z}$ from the graph-based radiative representation, defined as $V=\mathbf{V}(\mathcal{G})$. The voxelizer (Zha et al. 2024) divides the target space into $8\times 8\times 8$ 3D tiles and performs Gaussian culling to retain kernels with a 99% confidence of intersecting each tile. Building on this, we incorporate graph-based relationships into the voxelization process. The final voxel values are computed in parallel, as detailed in Eq. 13.
Optimization
To optimize the radiative Gaussians, we employ stochastic gradient descent. The total loss function consists of the photometric loss $\mathcal{L}_{1}$, the D-SSIM loss $\mathcal{L}_{ssim}$(Wang et al. 2004), the 3D total variation (TV) regularization $\mathcal{L}_{tv}$, and the graph Laplacian regularization $\mathcal{L}_{lap}$. The graph Laplacian regularization term $\mathcal{L}_{lap}$ is defined as:
$\mathcal{L}_{lap}(\mathcal{G})=\sum_{i=1}^{M}\sum_{j\in\mathcal{N}(i)}w_{ij}(\rho_{i}-\rho_{j})^{2},$ (18)
where $\mathcal{G}$ represents the graph structure of the Gaussian kernels, $\rho_{i}$ and $\rho_{j}$ denote the densities of Gaussian kernels $G_{i}^{3}$ and $G_{j}^{3}$, and $w_{ij}$ is the weight of the edge $(i,j)$. This term encourages local smoothness by minimizing density differences between neighboring kernels while preserving boundary information. To further enhance regularization, we define $\mathcal{L}_{norm}$ as:
$\mathcal{L}_{norm}=\lambda_{lap}\mathcal{L}_{lap}(\mathcal{G})+\lambda_{tv}\mathcal{L}_{tv}(\mathcal{V}_{tv}),$ (19)
where $\mathcal{L}_{tv}$ represents the 3D total variation regularization applied to the volume $\mathcal{V}_{tv}$. The total loss function is then expressed as:
$\mathcal{L}_{total}=\mathcal{L}_{1}(\mathbf{I}_{r},\mathbf{I}_{m})+\lambda_{ssim}\mathcal{L}_{ssim}(\mathbf{I}_{r},\mathbf{I}_{m})+\mathcal{L}_{norm}.$ (20)
Here, $\mathbf{I}_{r}$ and $\mathbf{I}_{m}$ denote the reconstructed and measured images, respectively.
Method HO BS AO Average RD
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
FDK 22.64 0.319 23.45 0.296 22.95 0.336 23.01 0.317 23.30 0.335
SART 29.58 0.773 32.37 0.878 31.48 0.825 31.14 0.825 31.52 0.790
ASD-POCS 29.42 0.810 31.40 0.887 30.58 0.845 30.47 0.847 31.32 0.810
NAF 32.05 0.841 34.36 0.930 35.34 0.909 33.92 0.893 32.92 0.772
SAX-NeRF 32.53 0.858 34.67 0.940 35.85 0.917 34.35 0.905 33.49 0.793
$R^{2}$-GS 32.98 0.881 35.08 0.944 37.52 0.945 35.19 0.922 35.03 0.837
Ours 33.47 0.891 35.56 0.952 38.55 0.955 35.86 0.933 35.95 0.858
Table 1: Detailed quantitative results under 25-view. Comparison of our GR-Gaussian with different methods on the X-3D dataset and Real-world dataset. HO: Human Organs, BS: Biological Specimens, AO: Artificial Objects (all from the X-3D dataset); RD: Real-world Dataset. We colorize the best, second-best, and third-best numbers.
5 Experiments
5.1 Experimental Settings
Dataset
Following $R^{2}$-Gaussian , we utilize the publicly available X-3D dataset to evaluate our method. This dataset(Zha et al. 2024) encompasses diverse categories, including human organs(chest, foot, head, jaw, and pancreas), artificial objects(backpack, engine, mount, present, and teapot), and biological specimens(beetle, bonsai, broccoli, kingsnake, and pepper), providing a comprehensive benchmark for reconstruction performance. Using the tomography toolbox TIGRE, we generate $512\times 512$ projections along a full circular trajectory, incorporating electric noise and ponton scatter to simulate realistic conditions. Subsequently, CT volumes are scanned within an angular range of $0^{\circ}$ to $360^{\circ}$ to produce 25-view projections. For the Real-World datasets(Society 2024), the projections are directly acquired using CT scanners and correspond to specific angles. Each projection has a resolution of $560\times 560$. To achieve sparse-view conditions, we uniformly select 25 projection angles from the full angular range of $0^{\circ}$ to $360^{\circ}$. This dataset including pine, seashell and walnut.
Implementation Details
The GR-Gaussian framework is implemented using PyTorch(Paszke 2019), PyTorch Geometric(Fey and Lenssen 2019), and CUDA(Guide 2013), and optimized with the Adam optimizer(Diederik 2014). Initial learning rates for location, density, scale, and rotation are set to 0.0002, 0.01, 0.005, and 0.001, respectively, with exponential decay reducing them to 10% of their initial values. Regularization parameters include a television volume level of $D=32$, loss weights $\lambda_{\mathrm{ssim}}=0.25$ and $\lambda_{\mathrm{tv}}=0.05$, a Graph Laplacian weight $\mathcal{L}_{\mathrm{lap}}=8\times 10^{-4}$, a Gaussian denoising parameter $\sigma_{d}=3$, and a gradient computation weight $\lambda_{\mathrm{g}}=1\times 10^{-4}$. The framework uses $M=50,000$ Gaussians, a density threshold $\tau=0.001$, and $k=6$ nearest neighbors. A dynamic stopping criterion ($Iter_{stop}$) evaluates PSNR every 500 iterations, terminating if PSNR decreases by more than 0.5%. All experiments were conducted on an RTX 4090 GPU. Quantitative evaluation employed PSNR and SSIM(Wang et al. 2004), with PSNR assessing 3D reconstruction accuracy and SSIM evaluating structural consistency across 2D slices.
Figure 5: The left shows CT reconstructions from the X-3D dataset, covering three categories: chest, bonsai, and teapot. The right displays Real-world dataset reconstructions with colorized slices to highlight details, all under 25-view conditions.
5.2 Results and Evaluation
The experiments aim to evaluate the performance of GR-Gaussian in sparse-view tomographic reconstruction, focusing on reconstruction quality across X-3D and Real-world datasets. On the X-3D dataset, GR-Gaussian outperforms the SOTA 3DGS-based method (Zha et al. 2024), NeRF-based methods (Zha, Zhang, and Li 2022; Cai et al. 2024b) and three traditional methods (Feldkamp, Davis, and Kress 1984; Andersen and Kak 1984; Sidky and Pan 2008) in both PSNR and SSIM. As show in Table 1, it achieves a PSNR increase of 0.67 dB and an SSIM improvement of 0.011, demonstrating its ability to suppress artifacts and enhance structural consistency. In the experiments with real-world datasets, GR-Gaussian demonstrates significant robustness against noise and artifacts, achieving a PSNR improvement of 0.92 dB and an SSIM improvement of 0.021. These results highlight the method’s adaptability to practical scenarios, where data imperfections pose significant challenges to reconstruction quality. As shown in Figure 5, GR-Gaussian produces visually superior reconstructions with sharper edges and fewer streak artifacts compared to other methods. These qualitative improvements align with the quantitative results, further validating the effectiveness of the proposed framework.
Baseline De-Init PGA X-3D Real-world
PSNR SSIM PSNR SSIM
35.20 0.923 35.28 0.836
35.61 0.931 35.90 0.857
35.77 0.932 35.62 0.848
35.86 0.932 35.95 0.858
Table 2: Ablation study of the components of the GR-Gaussian on the X-3D dataset and Real-world dataset.
Figure 6: Ablation study results highlight the impact of PGA and De-Init in enhancing reconstruction quality.
5.3 Ablation Study
Component Analysis
We performed ablation experiments to assess the contributions of Denoised Point Cloud Initialization (De-Init) and Pixel-Graph-Aware Gradient Strategy (PGA) to reconstruction performance on both the simulated X-3D dataset and the real-world dataset. For the ablation study, the baseline model omits the De-Init and PGA components, with Gaussians initialized via an FDK-based approach. To validate the component’s efficiency, we measure PSNR and SSIM under 25 views, with results listed in Table 2. Additionally, Figure 6 presents colorized volume and slice examples to visually highlight detailed improvements. Across the Real-world datasets, De-Init effectively mitigates needle-like artifacts, while PGA exhibits superior performance in preserving smooth regions.
k 4 5 6 7 8
X-3D PSNR 35.80 35.84 35.86 35.87 35.89
SSIM 0.932 0.932 0.933 0.933 0.934
Time 536s 567s 629s 763s 865s
Real World PSNR 35.82 35.87 35.95 35.97 36.01
SSIM 0.851 0.854 0.858 0.858 0.860
Time 562s 597s 669s 794s 927s
$\sigma_{d}$ 1 2 3 4 5
X-3D PSNR 35.60 35.72 35.86 35.68 35.43
SSIM 0.926 0.928 0.933 0.927 0.925
Real World PSNR 35.90 35.93 35.95 35.91 35.85
SSIM 0.857 0.858 0.858 0.857 0.856
Table 3: Performance comparison for different k-nearest neighbors and $\sigma_{d}$. Analysis of the X-3D dataset and Real-world dataset.
Parameter Analysis
The number of neighbors $k$ and the standard deviation $\sigma_{d}$ of the Gaussian kernel are critical hyperparameters in our method, influencing spatial relationship modeling and noise suppression. Sensitivity analysis on the X-3D and real-world datasets (Table 3) shows optimal reconstruction quality at $k=6$ and $\sigma_{d}=3$, achieving the highest PSNR and SSIM. Larger $k$ improves gradient accuracy by capturing richer neighborhoods but increases computational complexity, while smaller $k$ reduces overhead but compromises gradient precision. For De-Init, $\sigma_{d}$ directly affects filtering efficacy: smaller $\sigma_{d}$ preserves local details but inadequately suppresses noise, whereas larger $\sigma_{d}$ smooths noise effectively but risks over-smoothing edges and details.
Figure 7: Iteration Analysis Reconstruction results of GR-Gaussian across different iterations, illustrating the impact of iteration count on PSNR.
Iteration Analysis
Under sparse-view conditions, PSNR initially improves with increased iterations but eventually declines (see Fig. 7). This reduction is less pronounced in the simulated X-3D dataset, yet it becomes evident in real-world datasets, where noise and artifacts exacerbate reconstruction degradation. To address this issue, we employ a dynamic iteration stopping criterion ($Iter_{stop}$), where PSNR is evaluated every 500 iterations. If the PSNR decreases by more than 0.5%, the iteration process is terminated. This approach ensures that the model avoids overfitting and maintains optimal reconstruction quality.
6 Conclusion
This paper introduces GR-Gaussian, a novel framework based on 3D Gaussian Splatting for sparse-view tomographic reconstruction. Our approach employs a Denoised Point Cloud Initialization Strategy that robustly reduces noise while preserving critical structural details. Furthermore, by incorporating a graph-structured representation together with a novel gradient computation mechanism, GR-Gaussian significantly enhances reconstruction quality and effectively mitigates artifacts under challenging sparse-view conditions. Extensive evaluations on X-3D and real-world datasets demonstrate that our method outperforms existing techniques in both artifact suppression and reconstruction accuracy.

### v2

@@
 These results highlight the applicability of GR-Gaussian for accurate CT reconstruction under challenging sparse-view conditions.
-Figure 1: Results.
-We compare our method to two NeRF-based methods NAF(Zha, Zhang, and Li 2022), SAX-NeRF(Cai et al.
-2024b)) and $R^{2}$-GS(Zha et al.
-2024) in terms of visual quality and PSNR (dB).
+Fig 1: Visual results of GR-Gaussian.
+We compare GR-Gaussian to two NeRF-based methods NAF[44], SAX-NeRF[4]) and $R^{2}$-GS[45] in terms of visual quality and PSNR (dB).
 Our method mitigates needle-like artifacts and achieves superior CT reconstruction quality under sparse-view conditions.
+footnotetext: * Equal contribution.footnotetext: $\dagger$ Corresponding Authors.
 1 Introduction
-Computed Tomography (CT) is a widely used non-invasive imaging technique applied in fields such as medicine, biology, security, and industry.
-It utilizes X-rays to capture multi-angle 2D projections, which measure the attenuation of rays as they pass through an object(Kak and Slaney 2001).
-These projections are then processed to reconstruct the 3D density distribution of the object, enabling detailed internal visualization.
-In many applications, tomography data acquisition is limited by the number of projections or non-uniform angular sampling.
-For instance, medical CT reduces the number of projections to minimize X-ray exposure, while electron tomography faces the missing wedge problem due to restricted rotation angles.
-These constraints make the reconstruction problem highly ill-posed.
-Traditional analytical methods, such as FBP and FDK(Feldkamp, Davis, and Kress 1984), are computationally efficient but often fail under these conditions, producing artifacts.
-Model-based optimization approaches(Andersen and Kak 1984; Ma et al.
-2024b), which use nonlinear regularizers and iterative solvers (e.g., ART), achieve better results but are computationally expensive.
-Deep learning-based methods(Lin et al.
-2023; Chung et al.
-2023; Liu et al.
-2023) demonstrate strong performance but require large labeled datasets and long training times(Ma et al.
-2025a, c).
+Fig 2: Training pipeline of GR-Gaussian. (a) Overall training pipeline. (b) Denoised Point Cloud Initialization Strategy. (c) Pixel-Graph-Aware Gradient Strategy.
+ake the reconstruction problem highly ill-posed.
+Traditional analytical methods, such as FBP and FDK[8], are computationally efficient but often fail under these conditions, producing artifacts.
+Model-based optimization approaches[1, 28], which use nonlinear regularizers and iterative solvers (e.g., ART), achieve better results but are computationally expensive.
+Deep learning-based methods[23, 5, 24] demonstrate strong performance but require large labeled datasets and long training times[29, 31].
 They also struggle with out-of-distribution objects.
-NeRF-based(Mildenhall et al.
-2021; Ma et al.
-2025b, 2024a) approaches show promise in per-case reconstruction but are extremely time-consuming due to extensive point sampling during volume rendering(Zha, Zhang, and Li 2022; Cai et al.
-2024b; Zang et al.
-2021).
+NeRF-based[33, 30, 27] approaches show promise in per-case reconstruction but are extremely time-consuming due to extensive point sampling during volume rendering[44, 4, 43].
 These methods highlight the trade-off between reconstruction quality and computational efficiency.
-Recently, 3D Gaussian Splatting (3DGS)(Kerbl et al.
-2023; Yu et al.
-2024; Guédon and Lepetit 2024) has demonstrated superior quality and efficiency over NeRF in view synthesis tasks(Lu et al.
-2024; Liang et al.
-2024).
+Recently, 3D Gaussian Splatting (3DGS)[17, 41, 11] has demonstrated superior quality and efficiency over NeRF in view synthesis tasks[26, 21].
 However, when applied to sparse-view tomographic reconstruction, 3DGS-based methods exhibit characteristic needle-like artifacts.
@@
 Finally, we design an efficient optimization strategy tailored for GR-Gaussian, ensuring robust and accurate reconstruction.
-We validate our framework through extensive experiments on both simulated X-3D dataset(Zha et al.
-2024) and Real-world CT datasets (Society 2024), comparing it against multiple baseline methods.
+We validate our framework through extensive experiments on both simulated X-3D dataset[45] and Real-world CT datasets [38], comparing it against multiple baseline methods.
 The results demonstrate that GR-Gaussian outperforms existing approaches in reconstruction quality, achieving PSNR gains of 0.67 dB and 0.92 dB, and SSIM improvements of 0.011 and 0.021.
@@
   • Our framework achieves state-of-the-art performance on the X-3D and real-world datasets, demonstrating its effectiveness under challenging sparse-view conditions.
+Fig 3: The scanned object is represented as graph-based radiative Gaussians, optimized using real X-ray projections to retrieve the density volume via voxelization.
 2 Related work
 2.1 Tomographic Reconstruction
-Computed Tomography (CT) is a critical imaging technique widely used in medicine(Hounsfield 1980; Katsuragawa and Doi 2007) and industry(De Chiffre et al.
-2014; Zang et al.
-2019).
+Computed Tomography (CT) is a critical imaging technique widely used in medicine[14, 16] and industry[6, 42].
 Sparse-view CT reconstruction, where the number of projections is limited, poses significant challenges for traditional methods.
-Traditional analytical methods, such as Filtered Back Projection (FBP) and its 3D counterpart FDK(Feldkamp, Davis, and Kress 1984), are computationally efficient but prone to severe streak artifacts in sparse-view scenarios, leading to degraded image quality.
-Iterative methods(Andersen and Kak 1984; Sidky and Pan 2008; Manglos et al.
-1995; Sauer and Bouman 2002), which optimize reconstruction through regularized energy minimization, can suppress artifacts but are computationally expensive and risk losing fine structural details.
-Deep learning methods(Lin et al.
-2023; Chung et al.
-2023; Liu et al.
-2023) have shown promise in CT reconstruction, particularly in sparse-view scenarios.
-Supervised approaches leverage large datasets to learn semantic priors for tasks like projection inpainting(Anirudh et al.
-2018; Ghani and Karl 2018) and volume denoising(Chung et al.
-2023; Lee et al.
-2023; Liu et al.
-2023, 2020), but struggle with generalization to unseen data.
-Self-supervised methods, inspired by NeRF(Mildenhall et al.
-2021), optimize density fields using photometric losses(Cai et al.
-2024b; Zha, Zhang, and Li 2022; Zang et al.
-2021) but are computationally prohibitive due to extensive point sampling during rendering.
+Traditional analytical methods, such as Filtered Back Projection (FBP) and its 3D counterpart FDK[8], are computationally efficient but prone to severe streak artifacts in sparse-view scenarios, leading to degraded image quality.
+Iterative methods[1, 37, 32, 36], which optimize reconstruction through regularized energy minimization, can suppress artifacts but are computationally expensive and risk losing fine structural details.
+Deep learning methods[23, 5, 24] have shown promise in CT reconstruction, particularly in sparse-view scenarios.
+Supervised approaches leverage large datasets to learn semantic priors for tasks like projection inpainting[2, 10] and volume denoising[5, 18, 24, 25], but struggle with generalization to unseen data.
+Self-supervised methods, inspired by NeRF[33], optimize density fields using photometric losses[4, 44, 43] but are computationally prohibitive due to extensive point sampling during rendering.
 2.2 3D Gaussian Splatting
-3D Gaussian Splatting has demonstrated remarkable performance in RGB tasks, such as surface reconstruction(Li et al.
-2025; Lin et al.
-2025; Hou et al.
-2025), dynamic scene modeling(Zhu et al.
-2025; Huang et al.
-2025; Wu et al.
-2024), human avatar creation(Qu et al.
-2025; Li et al.
-2024), and 3D generation, outperforming NeRF in terms of speed by leveraging highly parallelized rasterization for image rendering.
+3D Gaussian Splatting has demonstrated remarkable performance in RGB tasks, such as surface reconstruction[19, 22, 13], dynamic scene modeling[46, 15, 40], human avatar creation[35, 20], and 3D generation, outperforming NeRF in terms of speed by leveraging highly parallelized rasterization for image rendering.
 Recent studies have attempted to extend 3DGS to X-ray imaging.
-For example, Cai(Cai et al.
-2024a) empirically modified 3DGS to synthesize novel-view X-ray projections, primarily using it as a data augmentation tool for traditional algorithms rather than directly generating CT models.
-Zha(Zha et al.
-2024) represented the density field with customized Gaussian kernels, yet they replaced the efficient rasterization with existing CT simulators.
+For example, Cai[3] empirically modified 3DGS to synthesize novel-view X-ray projections, primarily using it as a data augmentation tool for traditional algorithms rather than directly generating CT models.
+Zha[45] represented the density field with customized Gaussian kernels, yet they replaced the efficient rasterization with existing CT simulators.
 While effective in certain cases, their approach overfits sparse-view projections, leading to reconstruction failures.
-Figure 2: Training pipeline of GR-Gaussian. (a) Overall training pipeline. (b) Denoised Point Cloud Initialization Strategy. (c) Pixel-Graph-Aware Gradient Strategy.
-3 Preliminary
-3.1 Radiative Gaussians
-To represent the target object, previous approach(Zha et al.
-2024) utilize a set of learnable 3D Gaussian kernels, denoted as $\mathbb{G}^{3}=\{G_{i}^{3}\}_{i=1,\cdots,M}$, which we refer to as Radiative Gaussians.
+3 Method
+In this section, we present GR-Gaussian, a novel framework that represents objects through graph-structured relationships (Sec.
+3.2).
+Specifically optimized for sparse-view CT reconstruction, our framework effectively tackles critical issues, most notably the emergence of needle-like artifacts (Sec.
+3.3).
+We first introduce a Denoised Point Cloud Initialization Strategy based on an enhanced FDK method.
+Subsequently, a Pixel-Graph-Aware Gradient Strategy is proposed to refine the gradient computation process.
+Finally, we detail an optimization strategy specifically tailored for GR-Gaussian.
+3.1 Preliminaries
+3.1.1 Radiative Gaussians
+To represent the target object, previous approach[45] utilize a set of learnable 3D Gaussian kernels, denoted as $\mathbb{G}^{3}=\{G_{i}^{3}\}_{i=1,\cdots,M}$, which we refer to as Radiative Gaussians.
 Each kernel $G_{i}^{3}$ defines a local Gaussian-shaped density field, expressed as:
@@
 Notably, this formulation omits view-dependent color since X-ray attenuation depends solely on isotropic density, rendering it ideal for CT reconstruction.
-3.2 Pixel-Aware Gradient
+3.1.2 Pixel-Aware Gradient
 In 3D Gaussian Splatting, the decision to split or clone a Gaussian kernel $G_{i}^{3}$ is based on the average gradient magnitude of its Normalized Device Coordinates (NDC) across multiple viewpoints.
@@
 The exponential decay of $\alpha_{v,pix}^{i}$ emphasizes the influence of pixels closer to the Gaussian center, enabling precise gradient updates.
-4 Method
-In this section, we present GR-Gaussian, a novel framework that represents objects through graph-structured relationships (Sec.
-4.1).
-Specifically optimized for sparse-view CT reconstruction, our framework effectively tackles critical issues, most notably the emergence of needle-like artifacts (Sec.
-4.2).
-We first introduce a Denoised Point Cloud Initialization Strategy based on an enhanced FDK method.
-Subsequently, a Pixel-Graph-Aware Gradient Strategy is proposed to refine the gradient computation process.
-Finally, we detail an optimization strategy specifically tailored for GR-Gaussian.
-Figure 3: The scanned object is represented as graph-based radiative Gaussians, optimized using real X-ray projections to retrieve the density volume via voxelization.
 3.2 Representing Objects with Graph-based Radiative Gaussians
@@
 To ensure bidirectional connectivity, the edge set $\mathcal{E}$ is defined as:
-=$\mathcal{E}=\{(\mathbf{p}_{i},\mathbf{p}_{j})\mid\mathbf{p}_{j}\in\text{KNN}(\mathbf{p}_{i})\text{ and }\mathbf{p}_{i}\in\text{KNN}(\mathbf{p}_{j})\}.$ (9)
+$\mathcal{E}=\{(\mathbf{p}_{i},\mathbf{p}_{j})\mid\mathbf{p}_{j}\in\text{KNN}(\mathbf{p}_{i})\text{ and }\mathbf{p}_{i}\in\text{KNN}(\mathbf{p}_{j})\}.$ (9)
 Each Gaussian kernel $G_{i}^{3}$ defines a local density field by combining its own density with contributions from neighboring kernels:
@@
 Finally, a modified Pixel-Graph-Aware Gradient Strategy is applied to densify the Gaussians.
-Denoised Point Cloud Initialization Strategy
+3.3.1 Denoised Point Cloud Initialization Strategy
 Previous methods utilize FDK to generate low-quality volumes for initialization.
@@
 Empty regions are excluded using a density threshold $\tau$, and $M$ points are randomly sampled as kernel positions.
-Following (Zha et al.
-2024), Gaussian scales are set based on nearest neighbor distances, assuming no rotation.
+Following [45], Gaussian scales are set based on nearest neighbor distances, assuming no rotation.
 Central densities are directly queried from the denoised volume.
-Pixel-Graph-Aware Gradient Strategy
+3.3.2 Pixel-Graph-Aware Gradient Strategy
 Previous methods compute gradients based solely on average values, limiting the contribution of large kernels with small gradients to the densification process.
@@
 The Pixel-Graph-Aware Gradient Strategy integrates spatial relationships into gradient computation for 3D Gaussian Splatting, leveraging graph-based density differences to refine the splitting strategy.
-Figure 4: The Pixel-Graph-Aware Gradient Strategy leverages density differences between Gaussian kernels by constructing a graph to encode point-to-point relationships, enhancing gradient computation and enabling effective splitting of large kernels with low gradients.
-Density Voxelizer
+Fig 4: The Pixel-Graph-Aware Gradient Strategy leverages density differences between Gaussian kernels by constructing a graph to encode point-to-point relationships, enhancing gradient computation and enabling effective splitting of large kernels with low gradients.
+3.3.3 Density Voxelizer
 We employ a voxelizer, denoted as $\mathbf{V}$, to efficiently extract a density volume $V\in R^{X\times Y\times Z}$ from the graph-based radiative representation, defined as $V=\mathbf{V}(\mathcal{G})$.
-The voxelizer (Zha et al.
-2024) divides the target space into $8\times 8\times 8$ 3D tiles and performs Gaussian culling to retain kernels with a 99% confidence of intersecting each tile.
+The voxelizer [45] divides the target space into $8\times 8\times 8$ 3D tiles and performs Gaussian culling to retain kernels with a 99% confidence of intersecting each tile.
 Building on this, we incorporate graph-based relationships into the voxelization process.
@@
 13.
-Optimization
+3.3.4 Optimization
 To optimize the radiative Gaussians, we employ stochastic gradient descent.
-The total loss function consists of the photometric loss $\mathcal{L}_{1}$, the D-SSIM loss $\mathcal{L}_{ssim}$(Wang et al.
-2004), the 3D total variation (TV) regularization $\mathcal{L}_{tv}$, and the graph Laplacian regularization $\mathcal{L}_{lap}$.
+The total loss function consists of the photometric loss $\mathcal{L}_{1}$, the D-SSIM loss $\mathcal{L}_{ssim}$[39], the 3D total variation (TV) regularization $\mathcal{L}_{tv}$, and the graph Laplacian regularization $\mathcal{L}_{lap}$.
 The graph Laplacian regularization term $\mathcal{L}_{lap}$ is defined as:
@@
 HO: Human Organs, BS: Biological Specimens, AO: Artificial Objects (all from the X-3D dataset); RD: Real-world Dataset.
-We colorize the best, second-best, and third-best numbers.
 4 Experiments
 4.1 Experimental Settings
-Dataset
+4.1.1 Dataset
 Following $R^{2}$-Gaussian , we utilize the publicly available X-3D dataset to evaluate our method.
-This dataset(Zha et al.
-2024) encompasses diverse categories, including human organs(chest, foot, head, jaw, and pancreas), artificial objects(backpack, engine, mount, present, and teapot), and biological specimens(beetle, bonsai, broccoli, kingsnake, and pepper), providing a comprehensive benchmark for reconstruction performance.
+This dataset[45] encompasses diverse categories, including human organs(chest, foot, head, jaw, and pancreas), artificial objects(backpack, engine, mount, present, and teapot), and biological specimens(beetle, bonsai, broccoli, kingsnake, and pepper), providing a comprehensive benchmark for reconstruction performance.
 Using the tomography toolbox TIGRE, we generate $512\times 512$ projections along a full circular trajectory, incorporating electric noise and ponton scatter to simulate realistic conditions.
 Subsequently, CT volumes are scanned within an angular range of $0^{\circ}$ to $360^{\circ}$ to produce 25-view projections.
-For the Real-World datasets(Society 2024), the projections are directly acquired using CT scanners and correspond to specific angles.
+For the Real-World datasets[38], the projections are directly acquired using CT scanners and correspond to specific angles.
 Each projection has a resolution of $560\times 560$.
@@
 This dataset including pine, seashell and walnut.
-Implementation Details
-The GR-Gaussian framework is implemented using PyTorch(Paszke 2019), PyTorch Geometric(Fey and Lenssen 2019), and CUDA(Guide 2013), and optimized with the Adam optimizer(Diederik 2014).
+4.1.2 Implementation Details
+The GR-Gaussian framework is implemented using PyTorch[34], PyTorch Geometric[9], and CUDA[12], and optimized with the Adam optimizer[7].
 Initial learning rates for location, density, scale, and rotation are set to 0.0002, 0.01, 0.005, and 0.001, respectively, with exponential decay reducing them to 10% of their initial values.
@@
 All experiments were conducted on an RTX 4090 GPU.
-Quantitative evaluation employed PSNR and SSIM(Wang et al.
-2004), with PSNR assessing 3D reconstruction accuracy and SSIM evaluating structural consistency across 2D slices.
-Figure 5: The left shows CT reconstructions from the X-3D dataset, covering three categories: chest, bonsai, and teapot.
+Quantitative evaluation employed PSNR and SSIM[39], with PSNR assessing 3D reconstruction accuracy and SSIM evaluating structural consistency across 2D slices.
+Fig 5: The left shows CT reconstructions from the X-3D dataset, covering three categories: chest, bonsai, and teapot.
 The right displays Real-world dataset reconstructions with colorized slices to highlight details, all under 25-view conditions.
@@
 The experiments aim to evaluate the performance of GR-Gaussian in sparse-view tomographic reconstruction, focusing on reconstruction quality across X-3D and Real-world datasets.
-On the X-3D dataset, GR-Gaussian outperforms the SOTA 3DGS-based method (Zha et al.
-2024), NeRF-based methods (Zha, Zhang, and Li 2022; Cai et al.
-2024b) and three traditional methods (Feldkamp, Davis, and Kress 1984; Andersen and Kak 1984; Sidky and Pan 2008) in both PSNR and SSIM.
+On the X-3D dataset, GR-Gaussian outperforms the SOTA 3DGS-based method [45], NeRF-based methods [44, 4] and three traditional methods [8, 1, 37] in both PSNR and SSIM.
 As show in Table 1, it achieves a PSNR increase of 0.67 dB and an SSIM improvement of 0.011, demonstrating its ability to suppress artifacts and enhance structural consistency.
@@
 Table 2: Ablation study of the components of the GR-Gaussian on the X-3D dataset and Real-world dataset.
-Figure 6: Ablation study results highlight the impact of PGA and De-Init in enhancing reconstruction quality.
+Fig 6: Ablation study results highlight the impact of PGA and De-Init in enhancing reconstruction quality.
 4.3 Ablation Study
-Component Analysis
+4.3.1 Component Analysis
 We performed ablation experiments to assess the contributions of Denoised Point Cloud Initialization (De-Init) and Pixel-Graph-Aware Gradient Strategy (PGA) to reconstruction performance on both the simulated X-3D dataset and the real-world dataset.
@@
 Across the Real-world datasets, De-Init effectively mitigates needle-like artifacts, while PGA exhibits superior performance in preserving smooth regions.
-k 4 5 6 7 8
-X-3D PSNR 35.80 35.84 35.86 35.87 35.89
+Table 3: Quantitative results for different k-nearest neighbors and $\sigma_{d}$. Analysis of the X-3D dataset and Real-world dataset.
+Metric k=4 k=5 k=6 k=7 k=8
+X-3D Dataset
+PSNR 35.80 35.84 35.86 35.87 35.89
 SSIM 0.932 0.932 0.933 0.933 0.934
 Time 536s 567s 629s 763s 865s
-Real World PSNR 35.82 35.87 35.95 35.97 36.01
+Real World Dataset
+PSNR 35.82 35.87 35.95 35.97 36.01
 SSIM 0.851 0.854 0.858 0.858 0.860
@@
 $\sigma_{d}$ 1 2 3 4 5
-X-3D PSNR 35.60 35.72 35.86 35.68 35.43
+X-3D Dataset
+PSNR 35.60 35.72 35.86 35.68 35.43
 SSIM 0.926 0.928 0.933 0.927 0.925
-Real World PSNR 35.90 35.93 35.95 35.91 35.85
+Real World Dataset
+PSNR 35.90 35.93 35.95 35.91 35.85
 SSIM 0.857 0.858 0.858 0.857 0.856
-Table 3: Performance comparison for different k-nearest neighbors and $\sigma_{d}$. Analysis of the X-3D dataset and Real-world dataset.
-Parameter Analysis
+4.3.2 Parameter Analysis
 The number of neighbors $k$ and the standard deviation $\sigma_{d}$ of the Gaussian kernel are critical hyperparameters in our method, influencing spatial relationship modeling and noise suppression.
@@
 For De-Init, $\sigma_{d}$ directly affects filtering efficacy: smaller $\sigma_{d}$ preserves local details but inadequately suppresses noise, whereas larger $\sigma_{d}$ smooths noise effectively but risks over-smoothing edges and details.
-Figure 7: Iteration Analysis Reconstruction results of GR-Gaussian across different iterations, illustrating the impact of iteration count on PSNR.
-Iteration Analysis
+Table 4: Quantitative results of SSGU extension.
+Method w/o SSGU w/ SSGU $\lambda=3$ w/ SSGU $\lambda=5$
+SDS 40.37s 14.12s 8.62s
+DDS 66.89s 22.65s 13.90s
+CoSD 344.76s 128.97s 79.15s
+Fig 7: Iteration Analysis Reconstruction results of GR-Gaussian across different iterations, illustrating the impact of iteration count on PSNR.
+4.3.3 Iteration Analysis
 Under sparse-view conditions, PSNR initially improves with increased iterations but eventually declines (see Fig.